Aaron is 5 years older than Ryan. The sum of their ages is 27. How old are Aaron and Ryan?
Answer by Cromlix(4381)

Hi there,
Ryan = x
so,Aaron= x + 5
Ryan's age + Aaron's age = 27 years
x + (x + 5) = 27
2x + 5 = 27
2x + 5 - 5 = 27 - 5 Five off both sides)
2x = 22
x = 22/2
x = 11.
Ryan = 11 years old
Aaron = 16 years old
Hope this helps :-)